a faulty barometer reads 72.6cm mercury when the atmospheric pressure is 75.0cm mercury. calculate the atmospheric pressure when this barometer reads 72.0cm mercury
Answer #1

atmospheric pressure is 75.0cm and it reads 72.6 cm mercury.

so it reads 2.4 cm mercury for a 75 cm mercury atmospheric pressure.

so it will read 2.4/75 cm mecrury for a 1 cm of Atm pressure.

Hence it will reads (2.4/75)×72 cm mercury for a 72 cm mercury.

i.e ,( 2.4/75)*72 = 2.4cm mercury.

So atmopheric pressure on this value is,

P = 72cm mrcury + 2.4 cm mercury.

P = 74.4 cm mercury. ...Answer.,

2nd method..

by unitry method.

P = 75×72/72.6 = 74.4 cm mercury. ..Answer.

Hope this helps you.